# Evaluating LLMs’ Mathematical Reasoning in Financial Document Question Answering

### v2

@@ A.3 Categorization of LLM prompting errors in numerical tabular reasoning
Accr. Yield Carrying Amt. Allowance (Beginning) (Beginning) (Beginning) Beginning Balance \$27 \$920 \$95 Purchases \$1 \$130 \$0 Disposals/Payments Received \$-11 \$-594 \$0 Accretion \$-44 \$44 \$0 Builds (Reductions) to the Allowance \$0 \$0 \$-18 Increase to Expected Cash Flows \$-2 \$19 \$0 FX/Other \$17 \$-50 \$0 Balance at December 31, 2010 \$116 \$469 \$77 Question: What percentage did this total represent in relation to the carrying amount? Response #: Domain Knowledge: To find the percentage of a value (A) in relation to another value (B), divide the first value (A) by the second value (B) and multiply it by 100. To find the combined total, add the two values. Gold Evidences: The carrying amount of the company’s purchased distressed loan portfolio at December 31, 2010, was \$392 million. The net allowance was \$77 million as of December 31, 2010. Steps: 1. Find the combined total of net allowance and the carrying amount. Response 1: 392 + 77 = 469 2. Calculate the percentage of net allowance in relation to the carrying amount. Response 2: $\frac{77}{392}\times 100=19.64$ 3. Calculate the percentage of the combined total in relation to the carrying amount. Response 3: $\left(\frac{469}{392}\right)\times 100=119.64$ The final answer is 119.64%. Ground-truth steps: [(77/(392 + 77))*100] = 16.42% Reason: By the phrase in relation to the question expected a ratio of $quantity_{1}/quantity_{2}$ but the model returned $quantity_{2}/quantity_{1}$
@@
2001 2000 Gain on Sales of Depreciable Properties \$45428 \$52067 Gain on Land Sales \$5080 \$9165 Impairment Adjustment \$-4800 (4800) \$-540 (540) Total \$45708 \$60692 Question: What is the percent change in general and administrative expense from 2000 to 2001? Response #: Domain Knowledge: To find the percent change between two values, we need to find the difference between the new and old values, divide the difference by the old value, and multiply the result by 100. Gold Evidences: General and administrative expense decreased from \$21.1 million in 2000 to \$15.6 million for the year ended December 31, 2001. Steps: 1. Calculate the change in general and administrative expense by subtracting the 2001 value (\$15.6 million) from the 2000 value (\$21.1 million). Response 1: 21.1 - 15.6 = 5.5 2. Divide the difference obtained in Step 1 by the general and administrative expense in 2000 (\$21.1 million). Response 2: $\frac{5.5}{21.1}=0.2607$ 3. Multiply the result obtained in Step 2 by 100 to get the percent change. Response 3: $0.2607\times 100=26.07$ The final answer is 26.07%. Ground-truth steps: [((15.6 - 21.1)/21.31)*100] = -26.07% Reason: Incorrect Instantiation of variables resulting in a value with a flipped sign
@@
2013 2014 2015 2016 2017 CME Group Inc. \$164.01 \$194.06 \$208.95 \$279.85 \$370.32 S&P 500 \$132.39 \$150.51 \$152.59 \$170.84 \$208.14 Peer Group \$176.61 \$187.48 \$219.99 \$249.31 \$323.23 Question: What is the annualized return for s&p 500 from 2012 to 2017? Response #: Domain Knowledge: The annualized return refers to the average rate of return over a given period for an investment, which has been adjusted for compounding. To find the annualized return for the S&P 500 from 2012 to 2017, we need to use the formula: Annualized $\text{Annualized return}=\left(\left(\frac{\text{Ending value}}{\text{Startingvalue}}\right)^{\frac{1}{\text{Number of years}}}-1\right)$ Gold Evidences: Investment was \$100 on December 31, 2012;Value of S&P 500 investment on December 31, 2017 is \$208.14. Steps: 1. Divide the ending value (208.14) by the starting value (100) Response 1: $\frac{208.14}{100}=2.0814$ 2. Calculate the number of years elapsed (2017 - 2012) Response 2: 2017 - 2012 = 5 3. Raise the value obtained in #1 to the power of (1/number of years) Response 3: $(2.0814)^{\frac{1}{5}}=1.1606$ 4. Subtract 1 from the value obtained in #3 Response 4: 1.1606 - 1 = 0.1606 5. Multiply the value obtained in #4 by 100 to get the annualized return in percentage Response 5: 0.1606 * 100 = 16.06 The final answer is 16.06%. Ground-truth steps: [$(208.14/100)^{(}1/(2017-2012))$ 1] = 0.1577 * 100 = 15.77 Reason: Inaccurate calculation
@@
Figure 16: This illustration demonstrates how the "depth of hierarchy" is determined for each supporting piece of evidence. The depth of hierarchy corresponds to the number of levels or tiers of information structure within the table. The top level (1) encompasses the table itself, which contains information about the results of operations in a firm’s Advice & Wealth Management segment. The second level (2) includes columns specifying the years (2009, 2008, Change), creating a substructure within the table. The third level (3) involves the column "(in millions, except percentages)," and the fourth level (4) encompasses the rows under categories like "Revenues," "Distribution fees," "Net investment income," and so on. To locate specific values like "\$1733" and "\$1912", a retriever module needs to navigate through these four levels.

### v3

@@ A.4 Metadata Annotations Dataset Coverage
Accr. Yield Carrying Amt. Allowance (Beginning) (Beginning) (Beginning) Beginning Balance \$27 \$920 \$95 Purchases \$1 \$130 \$0 Disposals/Payments Received \$-11 \$-594 \$0 Accretion \$-44 \$44 \$0 Builds (Reductions) to the Allowance \$0 \$0 \$-18 Increase to Expected Cash Flows \$-2 \$19 \$0 FX/Other \$17 \$-50 \$0 Balance at December 31, 2010 \$116 \$469 \$77 Question: What percentage did this total represent in relation to the carrying amount? Response #: Domain Knowledge: To find the percentage of a value (A) in relation to another value (B), divide the first value (A) by the second value (B) and multiply it by 100. To find the combined total, add the two values. Gold Evidences: The carrying amount of the company’s purchased distressed loan portfolio at December 31, 2010, was \$392 million. The net allowance was \$77 million as of December 31, 2010.
Steps: 1. Find the combined total of net allowance and the carrying amount. Response 1: 392 + 77 = 469 2. Calculate the percentage of net allowance in relation to the carrying amount. Response 2: $\frac{77}{392}\times 100=19.64$ 3. Calculate the percentage of the combined total in relation to the carrying amount. Response 3: $\left(\frac{469}{392}\right)\times 100=119.64$ The final answer is 119.64%. Ground-truth steps: [(77/(392 + 77))*100] = 16.42% Reason: By the phrase in relation to the question expected a ratio of $quantity_{1}/quantity_{2}$ but the model returned $quantity_{2}/quantity_{1}$
@@
2001 2000 Gain on Sales of Depreciable Properties \$45428 \$52067 Gain on Land Sales \$5080 \$9165 Impairment Adjustment \$-4800 (4800) \$-540 (540) Total \$45708 \$60692 Question: What is the percent change in general and administrative expense from 2000 to 2001? Response #: Domain Knowledge: To find the percent change between two values, we need to find the difference between the new and old values, divide the difference by the old value, and multiply the result by 100. Gold Evidences: General and administrative expense decreased from \$21.1 million in 2000 to \$15.6 million for the year ended December 31, 2001.
Steps: 1. Calculate the change in general and administrative expense by subtracting the 2001 value (\$15.6 million) from the 2000 value (\$21.1 million). Response 1: 21.1 - 15.6 = 5.5 2. Divide the difference obtained in Step 1 by the general and administrative expense in 2000 (\$21.1 million). Response 2: $\frac{5.5}{21.1}=0.2607$ 3. Multiply the result obtained in Step 2 by 100 to get the percent change. Response 3: $0.2607\times 100=26.07$ The final answer is 26.07%. Ground-truth steps: [((15.6 - 21.1)/21.31)*100] = -26.07% Reason: Incorrect Instantiation of variables resulting in a value with a flipped sign
@@
2013 2014 2015 2016 2017 CME Group Inc. \$164.01 \$194.06 \$208.95 \$279.85 \$370.32 S&P 500 \$132.39 \$150.51 \$152.59 \$170.84 \$208.14 Peer Group \$176.61 \$187.48 \$219.99 \$249.31 \$323.23 Question: What is the annualized return for s&p 500 from 2012 to 2017? Response #: Domain Knowledge: The annualized return refers to the average rate of return over a given period for an investment, which has been adjusted for compounding. To find the annualized return for the S&P 500 from 2012 to 2017, we need to use the formula: $\text{Annualized return}=\left(\left(\frac{\text{Ending value}}{\text{Starting value}}\right)^{\frac{1}{\text{Number of years}}}-1\right)$ Gold Evidences: Investment was \$100 on December 31, 2012;Value of S&P 500 investment on December 31, 2017 is \$208.14.
Steps: 1. Divide the ending value (208.14) by the starting value (100) Response 1: $\frac{208.14}{100}=2.0814$ 2. Calculate the number of years elapsed (2017 - 2012) Response 2: 2017 - 2012 = 5 3. Raise the value obtained in #1 to the power of (1/number of years) Response 3: $(2.0814)^{\frac{1}{5}}=1.1606$ 4. Subtract 1 from the value obtained in #3 Response 4: 1.1606 - 1 = 0.1606 5. Multiply the value obtained in #4 by 100 to get the annualized return in percentage Response 5: 0.1606 * 100 = 16.06 The final answer is 16.06%. Ground-truth steps: $(208.14/100)^{(}1/(2017-2012))$ - 1] = 0.1577 * 100 = 15.77 Reason: Inaccurate calculation